### v1

# Optimal Hardness of Online Algorithms for Large Independent Sets

Abstract
We study the algorithmic problem of finding a large independent set in an Erdös-Rényi random graph $\mathbb{G}(n,p)$. For constant $p$ and $b=1/(1-p)$, the largest independent set has size $2\log_{b}n$, while a simple greedy algorithm—where vertices are revealed sequentially and the decision at any step depends only on previously seen vertices—finds an independent set of size $\log_{b}n$. In his seminal 1976 paper [Kar76], Karp challenged to either improve this guarantee or establish its hardness. Decades later, this problem remains open—one of the most prominent algorithmic problems in the theory of random graphs.
In this paper, we establish that a broad class of online algorithms, which we define, fails to find an independent set of size $(1+\varepsilon)\log_{b}n$ any constant $\varepsilon>0$, with high probability. This class includes Karp’s algorithm as a special case, and extends it by allowing the algorithm to query edges, not yet “seen” by the algorithm, which we call exceptional. Our lower bound holds throughout the regime $p\in[d/n,1-n^{-1/d}]$, where $d$ is a sufficiently large constant. In the dense regime, where $p$ is constant, we further prove that our result is asymptotically tight with respect to the number of exceptional edges queried, by designing an online algorithm which beats the half-optimality threshold when the number of exceptional edges is slightly larger than our bound.
Our result provides evidence for the algorithmic hardness of Karp’s problem by supporting the conjectured optimality of the aforementioned greedy algorithm and establishing it within the class of online algorithms. Our proof relies on a refined analysis of the geometric structure of tuples of large independent sets, establishing a variant of the Overlap Gap Property (OGP) commonly used as a barrier for classes of algorithms. While OGP has predominantly served as a barrier to stable algorithms, online algorithms are not stable. Thus, our application of OGP-based techniques to the online setting is novel.
1 Introduction
In this paper, we consider the algorithmic problem of finding a large independent set in the Erdös-Rényi random graph $\mathbb{G}(n,p)$, where each of the $\binom{n}{2}$ possible edges is present independently with probability $p:=p(n)$. For brevity, we set $b=b(p):=\frac{1}{1-p}$. Throughout the introduction, we focus on the dense regime where $p$ is constant in $n$, although our main result holds for a broader range of parameters.
Matula [Mat70, Mat76] observed that the largest independent set in $\mathbb{G}(n,p)$ is approximately of size $2\log_{b}n$ with high probability (w.h.p.), and stronger concentration results were established by Grimmett and McDiarmid [GM75] as well as Bollobás and Erdös [BE76]. Derived via the first and second moment method, these results are non-constructive. In his seminal 1976 paper, Karp [Kar76] observed that a simple greedy algorithm finds an independent set that is half-optimal—nearly of size $\log_{b}n$ w.h.p.—and raised the following question: devise a better algorithm finding an independent set of size $(1+\varepsilon)\log_{b}n$ for some $\varepsilon>0$ [Kar76, Open Problem 8]. Karp’s question is arguably one of the most prominent algorithmic problems in random graphs. Yet, decades later, it remains open. This highlights a gap between the existential result and the best known algorithmic guarantee—an early example of a statistical-computational gap—which motivates our work and prompts us to ask:
Can we design a better algorithm? If not, can we establish an algorithmic barrier?
The independent set problem in $\mathbb{G}(n,p)$ is perhaps the earliest known example exhibiting a statistical-computational gap. Despite extensive recent progress in understanding such gaps (see surveys [BPW18, Gam21, GMZ22, Gam25] and references below), no algorithmic improvement over Karp [Kar76] is known.
In this paper, we provide an evidence of the hardness of this problem by establishing its hardness within a class of online algorithms, which we define. This class includes Karp’s greedy algorithm as a special case. Recently, a similar type of hardness was established for the same problem within the class of algorithms based on low-degree polynomials [HS25]. We note however that the latter class does not include the greedy algorithm as a special case. A much earlier evidence of hardness was obtained in [Jer92] based on slow mixing of Markov Chain Monte Carlo (MCMC) methods for this problem. We note however that the failure of MCMC based methods is a questionable evidence of algorithmic hardness, since [GJX23] showed that in the related planted clique model, a modified variant of the MCMC can in fact recover planted cliques of size $\Omega(\sqrt{n})$ when initialized from the full graph.
1.1 Online Algorithms: Informal Description
We describe the class of online algorithms we consider informally and also compare it with similar algorithms considered in the past in the context of the same problem: finding independent sets in random graphs. As a motivation for the class of algorithms we consider, we focus on two essential features of the Karp’s greedy algorithm. First, observe that the algorithm decides the fate of the node $t$ at time $t$ by inspecting the connectivity of the subgraph induced by nodes $1,2,\ldots,t$. In other words, the information embedded in connectivity between nodes $t+1,\ldots,n$ and the remainder of the graph is not used in the decision process. This is one aspect of the “onlineness” of the algorithm: information being used at “time” $t$ is one that was accumulated by “time” $t$.
However, the greedy algorithm exhibits an online feature of a different kind as well. At time $t$ the algorithm decides on the fate of the node $t$ by inspecting its connectivity with previously constructed part $I_{t-1}$ of the independent set. Since $I_{t-1}$ has cardinality $O(\log n)$, the algorithm needs to inspect $O(\log n)$ edges at this step. The total number of inspections spanned by the independent set ultimately constructed is $O(\log^{2}n)$ (naturally, since the size is bounded as $O(\log n)$).
These two features are incorporated into the class of online algorithms we consider. Specifically, we consider the class of algorithms which make decisions about the fate of the node $t$ in step $t$ based on the following information: (a) inspecting all edges between nodes $u,v\in[t]$, and (b) inspecting a limited set of remaining ”future” edges $S_{t}$ in a restrictive way. In particular, the graph obtained from the intersection of $\cup_{t}S_{t}$ with the independent set ultimately constructed must have at most $c\log^{2}n$ edges. Our main result shows that online algorithms constrained in this way cannot improve upon the greedy algorithm when $c$ is small enough: for every $\varepsilon>0$ there exists $c=c(\varepsilon)$ such that the independent set produced by any algorithm satisfying the properties (a) and (b) is at most $(1+\varepsilon)\log_{b}n$ whp. We stress that this holds regardless of the algorithmic complexity of the algorithm considered. Its power is limited not by the time it is running but by the information it is using, and in this sense it is similar to the class of algorithms considered recently in [FGN+20, AHHM21, RS19, FF21], which we discuss below.
Perhaps surprisingly, our result is nearly tight, and furthermore, incorporating the restriction articulated in part (b) of the algorithmic feature is in fact well motivated. Specifically, we exhibit an algorithm which for every $\varepsilon\in(0,1)$ satisfies properties (a) and (b) with $c=3\varepsilon$, and ends up with an independent set of size $(1+\varepsilon)\log_{b}n$. See Theorem 2.4 for details.
Algorithms with limited look-ahead type features were considered recently starting from [FGN+20] and later on in [AHHM21, RS19, FF21] for the same problem of finding large independent sets in random graphs. In these papers a limited number of edges is inspected for the decision making process and, depending on the setting, independent sets of size $\alpha\log_{b}n$ with $\alpha>1$ can be constructed by algorithms. An essential difference between our setting and the setting considered in these papers is that in our case the fate of the node $t$ is decided in step $t$ and cannot be postponed to later stages. Contrasting with this, the aforementioned papers are based on a different “inspect first/decide later” type paradigm.
1.2 Statistical-Computational Gaps
Random optimization problems—optimization problems involving randomly generated inputs—are central to disciplines ranging from computer science and probability to statistics and beyond. Arising from a diverse array of models, many such problems interestingly share a common feature: a statistical-computational gap, namely a discrepancy between existential guarantees and what is algorithmically achievable. While the optimal value can often be identified using non-constructive techniques (e.g., the moment method), polynomial-time algorithms typically fall short of finding near-optimal solutions.
Models exhibiting such gaps go beyond the problem considered here and include random constraint satisfaction problems [MMZ05, ART06, ACO08, GS14, BH21], optimization over random graphs [GS14, RV17, GJW20, Wei20, PW24], binary perceptron and discrepancy minimization [GK23, GKPX22, GKPX23, Kız23a, LSZ24], spin glasses [CGPR19, GJ21, HS21, GJK23, Kız23b, HS23, Sel25], and more. For many of these models, all attempts at designing better algorithms have so far been unsuccessful, suggesting inherent algorithmic intractability.
Due to implicit randomness, standard worst-case complexity theory unfortunately offers little insight into such average-case hardness. [^1] For a few notable exceptions where formal hardness can be established under standard complexity-theoretic assumptions such as $P\neq\#P$, see, e.g., [Ajt96, BABB21, GK21, VV25]. Nevertheless, several powerful frameworks have emerged to understand such gaps and give evidence of computational hardness of random optimization problems, as well as other models, notably statistical problems involving planted structures. For a broader perspective on these approaches, see [WX18, BPW18, Gam21, GMZ22, Gam25].
1.3 Algorithmic Lower Bounds for Random Optimization Problems
For random optimization problems, one of the most powerful approaches to establishing algorithmic lower bounds leverages intricate geometry of the optimization landscape. While not explicitly ruling out algorithms, a body of work [MMZ05, ART06, ACO08] uncovered an intriguing connection for certain random computational problems: the onset of algorithmic hardness coincides with phase transitions in geometric features of the landscape. Gamarnik and Sudan [GS14] later initiated a formal framework known as Overlap Gap Property (OGP), which leverages these geometric properties to rigorously rule out broad classes of algorithms. [^2] The term OGP was formally introduced in [GL18]. Their focus was on the problem of finding a large independent set in sparse Erdös-Rényi random graph with average degree $d$, $\mathbb{G}(n,\frac{d}{n})$. It is known that the largest independent set in this graph has (w.h.p.) size $2\frac{\log d}{d}n$ in the double limit $n\to\infty$ followed by $d\to\infty$ [Fri90, FL92, BGT10]. In contrast, the best known algorithm finds an independent set of size only $\frac{\log d}{d}n$ [LW07], highlighting a factor-two gap mirroring our setting. Gamarnik and Sudan showed that independent sets of size $(1+1/\sqrt{2})\frac{\log d}{d}n$ either overlap substantially or are nearly disjoint, a structural property that leads to the failure of local algorithms. Later work by Rahman and Virag [RV17] established a sharp lower bound matching the algorithmic $\frac{\log d}{d}n$ value by considering the overlap pattern among many independent sets (multi-OGP). For the same algorithmic problem, lower bounds against low-degree polynomials—an emerging framework that capture implementations of many practical algorithms—were first obtained in [GJW20] and sharpened in [Wei20]. Since then, the OGP framework has found broad applicability beyond random graphs, including a variety of other random optimization problems.
For some problems, the algorithmic hardness predicted by the OGP-based technique actually matches a provable average case hardness. For example, a remarkable recent result in [VV25] shows that beating the threshold predicted by OGP for the binary perceptron model implies a polynomial-time algorithm for a worst-case instance of a variant of the shortest lattice vector problem, widely believed not solvable by polynomial time algorithms. At the same time some problems are in fact solvable by polynomial time algorithms beyond the OGP thresholds. In addition to the well known counterexample based on the random XOR-SAT instances which is solvable by Gaussian elimination, a more recent and interesting counterexample of this kind was constructed in [LS24] which shows that the shortest path problems, which is clearly solvable by polynomial-time algorithms, exhibits OGP in some random graph models. The shortest path problem can be solved by the linear programming technique and it is possible that problems admitting linear programming methods successfully overcome the OGP barrier. We do not attempt to review the (extensive) literature on OGP here, and instead refer the reader to several above mentioned surveys, in particular [Gam21, GMZ22, Gam25] and the references therein.
1.4 Prior Work: OGP as a Barrier to Stable Algorithms
[…]
Instability of Karp’s Algorithm
The success of the argument above hinges on algorithm’s stability, ensuring that it cannot “jump across” regions of forbidden overlap. As we demonstrate next, however, Karp’s greedy algorithm [Kar76] for finding an independent set of size $\log_{b}n$ is unstable, so it is unclear how to establish algorithmic lower bounds in this case.
Given a $G=(V,E)$ with $V=\{1,\dots,n\}:=[n]$ and any set $I$ of vertices, denote by $m(I)$ the smallest vertex not in $I$ and not adjacent to any $v\in I$. Initializing $I=\varnothing$, the algorithm in [Kar76] iteratively updates $S\leftarrow S\cup\{m(I)\}$ so long as $m(I)$ is defined. A nearly identical, though more accessible, version of this algorithm is due to Grimmett and McDiarmid [GM75]:
Algorithm 1 Greedy Algorithm for Independent Sets
  Input: A graph $G=(V,E)$ with $v=\{1,\dots,n\}$.
  Initialize: $I=\varnothing$
  for $i=1,\dots,n$ do  Set $I\leftarrow I\cup\{i\}$ if $i$ is not adjacent to any $j\in I$.
  end
  Output: $I$
A simple argument shows that for $G\sim\mathbb{G}(n,\frac{1}{2})$ and any $\varepsilon>0$, Algorithm 1 returns (w.h.p.) an independent set of size at least $(1-\varepsilon)\log_{2}n$ [GM75]. At the same time, we show:
Proposition 1.1 (Informal).
Algorithm 1 is unstable.
Proof Sketch.
Let $G_{1}\sim\mathbb{G}(n,\frac{1}{2})$ with edges $e_{1}(i,j)\sim{\rm Bern}(\frac{1}{2})$, $1\leq i<j\leq n$. Construct $G_{2}\sim\mathbb{G}(n,\frac{1}{2})$ by letting $e_{2}(i,j)=e_{1}(i,j)$ for all pairs except $\{i,j\}=\{1,2\}$, where $e_{2}(1,2)$ is resampled independently. Let $I_{1}$ and $I_{2}$ denote the outputs of running Algorithm 1 on $G_{1}$ and $G_{2}$, respectively. Observe that
$\mathbb{P}\bigl{[}v_{2}\notin I_{1},v_{2}\in I_{2}\bigr{]}\geq\mathbb{P}\bigl{[}e_{1}(1,2)=1,e_{2}(1,2)=0\bigr{]}=\frac{1}{4}.$
Moreover, $v_{2}$ has at least ${\rm Binomial}(|I_{1}|-1,\frac{1}{2})$ neighbors in $I_{1}$, which is at least $\frac{1-\varepsilon}{2}\log_{2}n$ (w.h.p.). Thus, on the event $\{v_{2}\notin I_{1},v_{2}\in I_{2}\}$—which holds w.p. $\Omega(1)$—none of these vertices are contained in $I_{2}$. Hence, simply by resampling one edge, $I_{1}$ and $I_{2}$ become substantially disjoint. ∎
Lower Bounds for Online Algorithms
Despite its instability, Algorithm 1 exhibits an online feature: decision at round $t$—i.e., whether $t\in I$—is based solely on the edges among the first $t$ vertices. While the OGP obstructs stable algorithms, geometric barriers beyond this class remain quite scarce. The idea of using OGP as a barrier to online algorithms was introduced in [GKPX23] for discrepancy minimization. The class of online algorithms considered there is rather straightforward and it is based on making decision based on the collected information without lookahead. This was well motivated though since the best known algorithm for discrepancy minimization constructed by Bansal and Spencer [BS20] is of this kind. Similar type of obstruction to online algorithms based on OGP was constructed recently for the graph alignment problem [DGH25]. Our class of online algorithms however is much richer and requires a far more delicate construction of the OGP based barrier. In particular, it necessitates defining a stopping time and considering multiple, carefully constructed interpolation paths.
1.5 Main Results (Informal)
Suppose that the vertices are probed in the order $v_{1},\dots,v_{t}$. In the simplest version, an algorithm $\mathcal{A}$ is online if the decision for $v_{t}$ is based solely on the information revealed up to round $t$—i.e., the edges among the first $t$ vertices.
Our focus is on a generalized class of online algorithms, which in particular captures Karp’s algorithm [Kar76] as a special case and goes as follows: at each step $t$, in addition to inspecting connectivity between the first $t$ nodes, the algorithm may additionally query an exceptional set of vertex pairs $S_{t}$, which may include pairs involving vertices not yet seen. The decision for vertex $t$ is then based both on (a) the edges among the first $t$ vertices, and (b) the edges among the exceptional pairs $E_{t}=S_{1}\cup\cdots\cup S_{t}$. For a formal statement, see Definition 2.1. In Section 2.2 we elaborate on exceptional pairs. We say that an algorithm $\mathcal{A}$ is $c$-good if
$\left|E_{n}\cap\binom{\mathcal{A}(G)}{2}\right|\leq c\log_{b}^{2}(np),$ (1)
where $\binom{\mathcal{A}(G)}{2}$ is the set of pairs $\{i,j\}$ with $i,j\in\mathcal{A}(G)$, and $\mathcal{A}(G)$ is the output of $\mathcal{A}$ on $G$ (see Definition 2.2). Note that Algorithm 1 trivially satisfies this condition with $c=0$.
Complementing Algorithm 1, our first main result establishes a sharp algorithmic lower bounds against general online algorithms.
Theorem 1.2 (Informal, see Theorem 2.3).
For any $\varepsilon>0$ and sufficiently small $c$, there exists no $c$-good algorithm finding an independent set of size $(1+\varepsilon)\log_{b}(np)$ w.p. at least $\exp\left(-\Omega\left(\log_{b}(np)\log(np)\right)\right)$.
Together with the guarantees of Karp [Kar76] and Grimmett and McDiarmid [GM75], Theorem 1.2 provides an additional evidence that $\log_{b}(np)$ is the algorithmic threshold for the problem of finding large independent sets in $\mathbb{G}(n,p)$. Furthermore, the probabilistic guarantee in Theorem 1.2 is essentially optimal, since the probability that a randomly chosen subset of size $(1+\varepsilon)\log_{b}(np)$ forms an independent set is $\exp\left(-O\left(\log_{b}(np)\log(np)\right)\right)$.
Our next main result concerns the sharpness of the bound (1) on the number of exceptional edges queried. In the dense regime (i.e., constant $p$), we show that condition (1) is tight up to constant factors:
Theorem 1.3 (Informal, see Theorem 2.4).
For any $\varepsilon\in(0,1)$, there exists a (deterministic) $3\varepsilon$-good algorithm that (w.h.p.) finds an independent set of size $(1+\varepsilon)\log_{b}(n)$.
Namely, the bound $O(\log_{b}^{2}(np))$ in (1) is essentially optimal up to constants. Of course, albeit being online, the algorithm achieving Theorem 1.3 runs in superpolynomial time. We also emphasize that for any $\varepsilon>0$, it is easy to construct a $3$-good algorithm—e.g., by brute-force searching for a large independent set and adding the vertices sequentially. Theorem 1.3 strengthens this observation by showing that for constant $p\in(0,1)$, one can in fact take $c\to 0$ as $\varepsilon\to 0$.
We emphasize that the online setting is far more than being just one algorithmic approach to solving a problem. It represents a fundamental computational model for decision-making under uncertainty, heavily studied in machine learning [RST10, RST11a, RST11b, RS13], convex optimization [H+16], and beyond. As mentioned, the best known algorithms for optimization over dense random graphs are online. Furthermore, for several important models—e.g., random k-SAT and combinatorial discrepancy theory [BS20]—the best known algorithms are either online or greedy, thus naturally admitting an online implementation. Given the central role that the online setting plays in modern data science, it is necessary to understand its fundamental limitations. Our results extend the existing toolkit and are potentially applicable to a broader range of problems.
2 Setup and Main Results
2.1 Online algorithms: Modus Operandi
Our focus is on the performance of following generalized class of online algorithms.
Definition 2.1.
Suppose that $G\sim\mathbb{G}(n,p)$ has vertex-set $[n]:=\{1,\dots,n\}$, whose edge-set is determined by the edge-status $e_{G}\bigl{(}\{u,v\}\bigr{)}\in\{0,1\}$ for all pairs $\{u,v\}\in\tbinom{[n]}{2}$. The randomized algorithm ${\mathcal{A}}$ whose internal randomness depend on seed $\omega$ proceeds in rounds, and at the end of each round $t\in[n]$ we shall keep track of the set $V_{t}=\{v_{1},\dots,v_{t}\}\subseteq[n]$ of vertices inspected by ${\mathcal{A}}$, the independent set $\mathcal{A}_{t}(G)\subseteq V_{t}$ produced so-far by ${\mathcal{A}}$, and the ‘exceptional’ pairs $E_{t}=S_{1}\cup\cdots\cup S_{t}\subseteq\binom{[n]}{2}$ inspected by ${\mathcal{A}}$, where initially $V_{0}:=\emptyset$, $\mathcal{A}_{0}(G):=\emptyset$, and $E_{t}=\emptyset$. Next, in each round $t\in[n]$ the algorithm ${\mathcal{A}}$ proceeds as follows.
  • (a)
    Based on $\omega$ and all the information revealed so far, the algorithm ${\mathcal{A}}$ selects a vertex $v_{t}\in[n]\setminus V_{t}$, and reveals the edge-status $e_{G}\bigl{(}\{v_{\ell},v_{t}\}\bigr{)}$ of all so-far unrevealed vertex-pairs $\{v_{\ell},v_{t}\}\in\tbinom{V_{t}}{2}$.
  • (b)
    Based on $\omega$ and all the information revealed so far, the algorithm ${\mathcal{A}}$ selects a set $S_{t}\subseteq\tbinom{[n]}{2}$ of vertex-pairs, and reveals the edge-status $e_{G}(\{u,v\})$ of all so-far unrevealed vertex-pairs $\{u,v\}\in S_{t}$.
  • (c)
    Based on $\omega$ and all the information revealed so far, the algorithm ${\mathcal{A}}$ decides whether ${\mathcal{A}_{t}(G)=\mathcal{A}_{t-1}(G)}$ or ${\mathcal{A}_{t}(G)=\mathcal{A}_{t-1}(G)\cup\{v_{t}\}}$, while maintaining that $\mathcal{A}_{t}(G)$ is an independent set of $G$.
At the end, we write $\mathcal{A}(G):=\mathcal{A}_{n}(G)$ for the independent set constructed by the algorithm ${\mathcal{A}}$.
$\mathcal{A}_{t}(G)$ is an independent set by construction for each $t\in[n]$. The sets $S_{t}$ are allowed to contain pairs $\{i,j\}$ with endpoints in $\{v_{t+1},\dots,v_{n}\}$—i.e., among the vertices yet to be explored. For this reason, we refer to edges with endpoints in $S_{t}$ as future edges. We also highlight the internal randomness of $\mathcal{A}$, captured by $\omega$, which affects the information revealed at any round.
2.2 Discussion of the Added Value of Exceptional Edges $S_{t}$
We next delve into the role of exceptional sets $S_{t}$. For simplicity, suppose that $\mathcal{A}$ probes the vertices in order $\{1,\dots,n\}$. Then, $S_{t}$ is allowed to contain pairs $\{i,j\}$ with $i\geq t+1$ or $j\geq t+1$, or both—that is, involving vertices that have not yet been probed by $\mathcal{A}$. Including $S_{t}$ makes the algorithm stronger by providing additional information that would otherwise be unavailable at round $t$. To see this, suppose that at round $t$, we observe $e(i,t)=0$ for all $i\in\mathcal{A}_{t-1}(G)$—i.e., the current vertex $t$ is not adjacent to any vertex in the independent set constructed thus far. Ignoring $S_{t}$, the natural greedy decision would be to add $t$ to the independent set:
$\mathcal{A}_{t}(G)\leftarrow\mathcal{A}_{t-1}(G)\cup\{t\}.$
Now, suppose that $S_{t}$ reveals a large subset of future vertices $j_{1},\dots,j_{K}\geq t+1$ such that:
  • (a)
    $e(i,j_{u})=0$ for all $i\in\mathcal{A}_{t-1}(G)$ and $\{i,j_{u}\}\in S_{t}$
  • (b)
    $e(t,j_{u})=1$ for all $u$.
In this case, including $t$ in the independent set—which may have seemed optimal without the extra information $S_{t}$—would prevent us to ever include $j_{1},\dots,j_{K}$ in future rounds. Thus, the ability to query exceptional pairs allows the algorithm to make more globally informed decisions.
Condition (1) and the bound $c\log_{b}^{2}(np)$ controls the total number of such exceptional edges. In what follows, we argue the sharpness of $\log_{b}^{2}(np)$ scaling by showing that: (a) for any $\varepsilon>0$ and $c$ sufficiently small, online algorithms fail to identify an independent set of size $(1+\varepsilon)\log_{b}(np)$; whereas (b) there exists an online algorithm with $c=3\varepsilon$ that finds an independent set of this size.
2.3 Formal Definition and Main Result
Note that besides $G$, $\mathcal{A}$ has additional randomness: (a) the order at which it explores the vertices, and (b) the exceptional sets $S_{t}$ revealed at any round $t$. To capture this, suppose there exists a probability space $(\Omega,\mathbb{P}_{\omega})$ such that
$\mathcal{A}:\{0,1\}^{\binom{n}{2}}\times\Omega\to\{0,1\}^{n}.$
We next define the $c$-good algorithms:
Definition 2.2.
Given $c>0$, we say that an algorithm $\mathcal{A}$ is $c$-good if:
$\left|\binom{\mathcal{A}_{n}(G)}{2}\cap E_{n}\right|\leq c\log^{2}_{b}(np).$
We now state our main result formally.
Theorem 2.3 (Main result).
Given $\varepsilon>0$, there exist constants $c,\xi>0$ and $d,n_{0}>1$ (which may depend on $\varepsilon)$ such that, for all $n\geq n_{0}$, any $c$-good algorithm $\mathcal{A}$ satisfies
$\mathbb{P}_{G\sim\mathbb{G}(n,p),\omega}\bigl{[}\bigl{|}\mathcal{A}(G,\omega)\bigr{|}\geq(1+\varepsilon)\log_{b}(np)\bigr{]}\leq(np)^{-\xi\log_{b}(np)}$ (2)
for all edge-probabilities $p=p(n)$ satisfying $d/n\leq p\leq 1-n^{-1/d}$.
Our proof uses $c=\min\{\varepsilon^{2},1\}/8$ and $\xi=\min\{\varepsilon^{2},1\}/64$. Furthermore, in (2), we have:
$(np)^{-\xi\log_{b}(np)}\leq n^{-\Omega(1)}\to 0$
as $n\to\infty$, as $\log_{b}(np)=\Omega(\log d/p)$ by the estimate (5) below. We prove Theorem 2.3 in Sections 3 and 4.
2.4 Sharpness of Definition 2.2: an Algorithmic Guarantee
Focusing on the dense regime where $p$ is constant, we now switch to our second result regarding the sharpness of our bound on $c$-goodness in Definition 2.2. Theorem 2.4 asserts that our algorithmic lower bound in Theorem 2.3 is nearly tight with respect to number of exceptional edges. Furthermore, it underscores that Part (b) of Definition 2.1 is indeed well motivated, and that $O(\log^{2}n)$ is the “right scaling”: there is a $3\varepsilon$-good online algorithm that can identify an independent set of size $(1+\varepsilon)\log_{b}n$.
Theorem 2.4.
For any fixed $\varepsilon,p\in(0,1)$, there exists a deterministic $3\varepsilon$-good algorithm $\mathcal{A}$ that satisfies
$\mathbb{P}_{G\sim\mathbb{G}(n,p)}\bigl{[}\bigl{|}\mathcal{A}(G)\bigr{|}\geq(1+\varepsilon)\log_{b}(n)\bigr{]}\to 1$ (3)
as $n\to\infty$.
This algorithm has running time $n^{(2+o(1))\varepsilon\log_{b}(n)}$ which is (naturally) superpolynomial. The proof is found in Section 5 for further details.
3 Impossibility Result: Proof of Theorem 2.3
In this section, we prove our main impossibility result Theorem 2.3, which shows that certain online algorithms cannot find independent sets of size $(1+\varepsilon)\log_{b}(np)$ with non-negligible probability.
To this end we may without loss of generality assume that $\varepsilon\leq 1$ (since the desired estimate (2) for $\varepsilon\geq 1$ is implied by estimate (2) for $\varepsilon=1$). Given $\varepsilon\in(0,1]$, with foresight we define
$c:=\frac{\varepsilon^{2}}{8},\qquad\xi:=\frac{\varepsilon^{2}}{64},\quad\text{and }\quad m:=\left\lceil\frac{16}{\varepsilon^{2}}\right\rceil,$ (4)
and defer the choice of the constants $d,n_{0}>1$. We henceforth fix a $c$-good algorithm $\mathcal{A}$ per Definition 2.2.
To avoid clutter, we henceforth always tacitly assume that $n$ is sufficiently large whenever necessary (how large may depend on all other parameters $\varepsilon,c,\xi,m,d$).
The parameter $\log_{b}(np)$ is a bit cumbersome to work with, since in different parts of the edge-probability range $d/n\leq p\leq 1-n^{-1/d}$, it behaves quite differently. For this reason, we now record a convenient lower bound based on standard asymptotic estimates:
Lemma 3.1.
There exists a universal constant $c_{0}>0$ (which does not depend on $\varepsilon,c,\xi,m,d,n_{0}$) such that
$\log_{b}(np)=\frac{\log(np)}{-\log(1-p)}\geq\frac{c_{0}\log d}{p}.$ (5)
Proof.
Note that $-\log(1-p)=p+O(p^{2})$ for $p\leq 1/2$, say. Choosing $20<p_{0}\leq 1/$ 2 sufficiently small, we thus obtain that for $d/n\leq p\leq p_{0}$, we have $-\log(1-p)\leq 9p$ and $\log(np)\geq\log d$. Consequently, $\log_{b}(np)\geq(\log d)/(9p)$ in (5).
In the remaining case, $p_{0}\leq p\leq 1-n^{-1/d}$, the crux is that $-\log(1-p)\leq\log(n)/d$ and $\log(np)\geq\log(np_{0})\geq\log(n)/2$ for sufficiently large $n$ (depending only on $p_{0}$), so that $\log_{b}(np)\geq d/2\geq p_{0}d/(2p)$ in (5). This establishes (5) by checking that $\min\{\log d,\>d\}=\log d$ for all $d\geq 1$. ∎
3.1 Reduction to deterministic algorithm
Note that the algorithm $\mathcal{A}$ in Theorem 2.3 is randomized: we shall use a standard averaging-based reduction, to show that in our upcoming arguments it suffices to consider a deterministic algorithm.
Lemma 3.2.
There exists ${\omega^{*}\in\Omega}$ such that the deterministic algorithm ${\mathcal{A}(G):=\mathcal{A}(G,\omega^{*})}$ satisfies
$\mathbb{P}_{G\sim\mathbb{G}(n,p),\omega}\bigl{[}|\mathcal{A}(G,\omega)|\geq(1+\varepsilon)\log_{b}(np)\bigr{]}\leq\mathbb{P}_{G\sim\mathbb{G}(n,p)}\bigl{[}|\mathcal{A}(G)|\geq(1+\varepsilon)\log_{b}(np)\bigr{]}.$ (6)
Proof.
By explicitly averaging over the randomness $\omega\in\Omega$, we can write
$\mathbb{P}_{G\sim\mathbb{G}(n,p),\omega}\bigl{[}|\mathcal{A}(G,\omega)|\geq(1+\varepsilon)\log_{b}(np)\bigr{]}=\mathbb{E}_{\omega}\Bigl{[}\mathbb{P}_{G\sim\mathbb{G}(n,p)}\bigl{[}|\mathcal{A}(G,\omega)|\geq(1+\varepsilon)\log_{b}(np)\bigr{]}\Bigr{]}.$
Hence there exists a choice of ${\omega^{*}\in\Omega}$ such that ${\mathcal{A}(G)=\mathcal{A}(G,\omega^{*})}$ satisfies estimate (6). ∎
We henceforth fix a choice $\omega^{*}\in\Omega$ that is guaranteed by Lemma 3.2, and set $\mathcal{A}(G):=\mathcal{A}(G,\omega^{*})$. In the remainder of the proof of Theorem 2.3, our goal will be to derive an upper bound on the success probability of the deterministic algorithm $\mathcal{A}$ (which suffices by estimate (6) from Lemma 3.2). Conceptually, the main advantage of working with a deterministic algorithm ${\mathcal{A}}$ is that the only source of randomness is the graph on which ${\mathcal{A}}$ operates.
3.2 Interpolation paths: correlated random graphs
The strategy of our proof of Theorem 2.3 is to analyze the behavior of the algorithm ${\mathcal{A}}$ on multiple correlated random graph families, referred to as interpolation paths. To this end we define
$G_{1}\sim\mathbb{G}(n,p)$ (7)
as an independent copy of the binomial random graph $\mathbb{G}(n,p)$, i.e., where each possible pair of vertices $\{u,v\}\in\tbinom{[n]}{2}$ is included as an edge independently with probability $p$. We shall run the algorithm ${\mathcal{A}}$ on $G_{1}$ to dynamically construct the families
$G_{1}^{(T)},\dots,G_{m}^{(T)}\qquad\text{for $1\leq T\leq n$}$
of correlated random graphs as follows, where to avoid clutter we henceforth write
$V_{T}=V_{T}(G_{1})=\{v_{1},\ldots,v_{T}\}\qquad\text{ and }\qquad E_{T}=E_{T}(G_{1})=S_{1}\cup\cdots\cup S_{T}.$ (8)
For the first graph sequence $G_{1}^{(T)}$ we simply set
$G^{(T)}_{1}:=G_{1}\qquad\text{for all~{}$1\leq T\leq n$,}$ (9)
and the basic idea for the other graphs $G_{i}^{(T)}$ with $1\leq T\leq n$ is heuristically speaking as follows: the edge-status of all pairs revealed by ${\mathcal{A}}$ during the first $T$ rounds is the same in $G_{1}$ and $G_{i}^{(T)}$, and all other pairs are included as an edge into $G_{i}^{(T)}$ independently with probability $p$. More formally, denoting by $e_{i,T}\bigl{(}\{u,v\}\bigr{)}\in\{0,1\}$ the edge-status of the pair $\{u,v\}$ in $G_{i}^{(T)}$, we define
$e_{i,T}\bigl{(}\{u,v\}\bigr{)}:=e_{1,T}\bigl{(}\{u,v\}\bigr{)}\qquad$ for all $1\leq T\leq n$, $2\leq i\leq T$ and $\{u,v\}\in\tbinom{V_{T}}{2}\cup E_{T}$, (10)
$e_{i,T}\bigl{(}\{u,v\}\bigr{)}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}{\rm Bernoulli}(p)\qquad$ for all $1\leq T\leq n$, $2\leq i\leq T$ and $\{u,v\}\not\in\tbinom{V_{T}}{{}_{2}}\cup E_{T}$. (11)
We now record some basic properties of the correlated random graphs $G_{1}^{(T)},\dots,G_{m}^{(T)}$.
Remark 3.3.
For any $1\leq i\leq m$ and $1\leq T\leq n$ we have $G_{i}^{(T)}\sim\mathbb{G}(n,p)$, since definitions (9) and (10)–(11) ensure that each potential edge of $G^{(T)}_{i}$ is included independently with probability $p$.
Remark 3.4.
For any $1\leq T\leq n$, the first $T$ rounds of the algorithm ${\mathcal{A}}$ on each graph $G_{1}^{(T)},\dots,G_{m}^{(T)}$ behave exactly the same as on the graph $G_{1}$, since definitions (10) and (9) ensure that all pairs revealed during those rounds have exactly the same edge-status as in $G^{(T)}_{1}=G_{1}$ (which by Definition 2.1 is all the information that the deterministic algorithm $\mathcal{A}$ uses to makes it selections and decisions).
3.3 Comparison argument via successful event ${\mathcal{S}}$
Our proof of Theorem 2.3 hinges on the careful definition of the ‘successful’ event ${\mathcal{S}}$ in (13) below, which captures the idea that the algorithm ${\mathcal{A}}$ finds an independent set of size at least $(1+\varepsilon)\log_{b}(np)$ in the correlated random graphs from Section 3.2. Turning to the details, setting with foresight
$\tau:=\min\biggl{\{}\min\Bigl{\{}1\leq t\leq n:\>|\mathcal{A}_{t}(G_{1})|=\left\lceil(1-\varepsilon/2)\log_{b}(np)\right\rceil\Bigr{\}},\ n\biggr{\}},$ (12)
we define the ‘successful’ event
${\mathcal{S}}:=\bigcap_{1\leq i\leq m}\Bigl{\{}\bigl{|}{\mathcal{A}}(G_{i}^{(\tau)})\bigr{|}\geq(1+\varepsilon)\log_{b}(np)\Bigr{\}}=\bigcap_{0\leq i\leq m}{\mathcal{E}}_{i,\tau},$ (13)
where ${\mathcal{E}}_{i,T}$ denotes the event that ${\bigl{|}{\mathcal{A}}(G_{i}^{(T)})\bigr{|}}\geq{(1+\varepsilon)\log_{b}(np)}$. The strategy of our comparison-based argument is to estimate the probability $\mathbb{P}[{\mathcal{S}}]$ in two different ways (from below and above), which eventually establishes the desired probability bound (2) of our main result Theorem 2.3; see Section 3.4.
Remark 3.5.
Definitions (12) and (13) of $\tau$ and ${\mathcal{S}}$ are an important ingredient of our proof. Specifically, in the proof of Lemma 3.6 below they are key for avoiding the need to take a union bound over all possible $1\leq T\leq n$ where $|\mathcal{A}_{T}(G_{1})_{T}|=\left\lceil(1-\varepsilon/2)\log_{b}(np)\right\rceil$ might occur, which in turn is crucial for being able to show that the resulting probability bound (2) goes to zero in our main result Theorem 2.3.
3.3.1 Lower bound on $\mathbb{P}[{\mathcal{S}}]$: conditioning and convexity
The following lemma estimates the ‘success’ probability $\mathbb{P}[{\mathcal{S}}]$ from below, by relating it to the probability $\mathbb{P}[{\mathcal{E}}]$ that the algorithm ${\mathcal{A}}$ finds an independent set of size at least $(1+\varepsilon)\log_{b}(np)$ in $G_{1}\sim\mathbb{G}(n,p)$. Concretely, the below proof of (14) combines the careful definitions (12) and (13) of $\tau$ and ${\mathcal{S}}$ with an application of Jensen’s inequality and the fact that, after suitable conditioning, certain events become conditionally independent (due to the way the algorithm ${\mathcal{A}}$ makes its decisions, see Definition 2.1).
Lemma 3.6.
Denoting by ${\mathcal{E}}$ the event that ${\bigl{|}{\mathcal{A}}(G_{1})\bigr{|}}\geq{(1+\varepsilon)\log_{b}(np)}$, we have
$\mathbb{P}[{\mathcal{S}}]\geq\mathbb{P}[{\mathcal{E}}]^{m}.$ (14)
Proof.
Let $\Xi_{T}$ encode the edge-status of all pairs $\{u,v\}\in\tbinom{V_{T}}{2}\cup E_{T}$, i.e., the edge-status of all pairs revealed during the first $T$ rounds of the algorithm ${\mathcal{A}}$ (which intuitively corresponds to the common randomness of $G^{(T)}_{1},\ldots,G^{(T)}_{m}$). We have
$\begin{split}\mathbb{P}[\tau=T,\;{\mathcal{S}}]&=\mathbb{P}[\tau=T,\;{\mathcal{E}}_{1,T},\dots,{\mathcal{E}}_{m,T}]\\ &=\mathbb{E}\bigl{[}\mathbb{P}[\tau=T,\;{\mathcal{E}}_{1,T},\dots,{\mathcal{E}}_{m,T}\>|\>\Xi_{T}]\bigr{]}\\ &=\mathbb{E}\bigl{[}\mathbbm{1}\{\tau=T\}\mathbb{P}[{\mathcal{E}}_{1,T},\dots,{\mathcal{E}}_{m,T}\>|\>\Xi_{T}]\bigr{]},\end{split}$ (15)
where the last equality in (15) follows from the fact that the event $\tau=T$ is determined by $\Xi_{T}$. Recall that ${\mathcal{E}}_{i,T}$ denotes the event that ${\bigl{|}{\mathcal{A}}(G_{i}^{(T)})\bigr{|}}\geq{(1+\varepsilon)\log_{b}(np)}$, where $\mathcal{A}$ is a deterministic algorithm that runs on the random graph $G_{i}^{(T)}$. Note that by (11), conditional on the edge-status $\Xi_{T}$ of the first $T$ rounds, all remaining pairs $\{u,v\}\not\in\tbinom{V_{T}}{2}\cup E_{T}$ are included as edges into each graph $G_{1}^{(T)},\ldots,G_{m}^{(T)}$ independently with probability $p$. Combined with Remark 3.4 and $G_{1}^{(T)}=G_{1}$, it thus follows that
$\mathbb{P}[{\mathcal{E}}_{1,T},\dots,{\mathcal{E}}_{m,T}\>|\>\Xi_{T}]$ $=\prod_{1\leq i\leq m}\mathbb{P}[{\mathcal{E}}_{i,T}\>\big{|}\>\Xi_{T}]=\mathbb{P}[{\mathcal{E}}_{1,T}\mid\Xi_{T}]^{m}=\mathbb{P}[{\mathcal{E}}\mid\Xi_{T}]^{m}.$ (16)
Since $1\leq\tau\leq n$ holds deterministically by definition (12), after inserting (16) into (15) it follows that
$\begin{split}\mathbb{P}[{\mathcal{S}}]&=\sum_{1\leq T\leq n}\mathbb{P}[\tau=T,\;{\mathcal{S}}]\\ &=\sum_{1\leq T\leq n}\mathbb{E}\bigl{[}\mathbbm{1}\{\tau=T\}\mathbb{P}[{\mathcal{E}}\mid\Xi_{T}]^{m}\bigr{]}\\ &=\mathbb{E}\Bigl{[}\sum_{1\leq T\leq n}\mathbbm{1}\{\tau=T\}\mathbb{P}[{\mathcal{E}}\mid\Xi_{T}]^{m}\Bigr{]}.\end{split}$ (17)
We note that conditionally on $\Xi_{1},\ldots,\Xi_{n}$ we have
$\sum_{1\leq T\leq n}\mathbbm{1}\{\tau=T\}\mathbb{P}[{\mathcal{E}}\mid\Xi_{T}]^{m}=\Bigl{(}\sum_{1\leq T\leq n}\mathbbm{1}\{\tau=T\}\mathbb{P}[{\mathcal{E}}\mid\Xi_{T}]\Bigr{)}^{m},$
which, by taking expectations, implies together with (17) that
$\mathbb{P}[{\mathcal{S}}]=\mathbb{E}\biggl{[}\Bigl{(}\sum_{1\leq T\leq n}\mathbbm{1}_{\{{\tau=T}\}}\mathbb{P}[{\mathcal{E}}\mid\Xi_{T}]\Bigr{)}^{m}\biggr{]}.$ (18)
Using that $x\mapsto x^{m}$ is convex for every $x\geq 0$ (due to $m\geq 1$), by applying Jensen’s inequality to the right-hand side of (18) it follows that
$\begin{split}\mathbb{P}[{\mathcal{S}}]\geq\biggl{(}\mathbb{E}\Bigl{[}\sum_{1\leq T\leq n}\mathbbm{1}_{\{{\tau=T}\}}\mathbb{P}[{\mathcal{E}}\mid\Xi_{T}]\Bigr{]}\biggr{)}^{m}.\end{split}$ (19)
Since $1\leq\tau\leq n$ holds deterministically by definition (12), we also have
$\begin{split}\mathbb{E}\Bigl{[}\sum_{1\leq T\leq n}\mathbbm{1}\{\tau=T\}\mathbb{P}[{\mathcal{E}}\mid\Xi_{T}]\Bigr{]}=\sum_{1\leq T\leq n}\mathbb{E}\bigl{[}\mathbbm{1}\{\tau=T\}\mathbb{P}[{\mathcal{E}}\mid\Xi_{T}]\bigr{]}=\sum_{1\leq T\leq n}\mathbb{P}[\tau=T,\;{\mathcal{E}}]=\mathbb{P}[{\mathcal{E}}],\end{split}$ (20)
where the fact that the event $\tau=T$ is determined by $\Xi_{T}$ implies
$\mathbb{E}\bigl{[}\mathbbm{1}\{\tau=T\}\mathbb{P}[{\mathcal{E}}\mid\Xi_{T}]\bigr{]}=\mathbb{E}\bigl{[}\mathbb{P}[\tau=T,\;{\mathcal{E}}|\Xi_{T}]\bigr{]}=\mathbb{P}[\tau=T,\;{\mathcal{E}}].$
Finally, combining (19) with (20) completes the proof of estimate (14) from Lemma 3.6. ∎
3.3.2 Upper bound on $\mathbb{P}[{\mathcal{S}}]$: unlikely independent sets
To estimate the ‘success’ probability $\mathbb{P}[{\mathcal{S}}]$ defined in Section 3.3 from above, the idea is to show that the event ${\mathcal{S}}$ implies the existence of a certain specific family of independent sets, which intuitively only exists with very small probability. The following definition and the subsequent lemmas formalize this idea, where $V_{T}=\{v_{1},\ldots,v_{T}\}$ denotes the first $T$ vertices considered by the algorithm, and $E_{T}=S_{1}\cup\cdots\cup S_{T}$ contains the ‘exceptional’ pairs revealed in (b) during the first $T$ rounds of the algorithm.
Definition 3.7.
Let $N:=\left\lceil(1-\varepsilon/2)\log_{b}(np)\right\rceil$. For any integer $1\leq T\leq n$, we denote by $X_{m,T}$ the number of $m$-tuples of sets $(I_{1},\dots,I_{m})$ with $I_{1},\dots,I_{m}\subseteq[n]$ that satisfy the following three properties:
  1. (i)
    for every $1\leq i\leq m$, the set $I_{i}$ is an independent set in $G_{i}^{(T)}$, with $|I_{i}|\geq(1+\varepsilon)\log_{b}(np)$,
  2. (ii)
    we have ${I_{i}\cap V_{T}}={I_{1}\cap V_{T}}$ for every $1\leq i\leq m$, with $|I_{1}\cap V_{T}|=N$, and
  3. (iii)
    we have $\bigl{|}\binom{I_{i}}{2}\cap E_{T}\bigr{|}\leq c\log_{b}^{2}(np)$ for every $1\leq i\leq m$.
Lemma 3.8.
We have
$\mathbb{P}[{\mathcal{S}}]\leq\sum_{1\leq T\leq n}\mathbb{P}[X_{m,T}\geq 1].$ (21)
Proof.
Since $1\leq\tau\leq n$ holds deterministically by definition (12), it follows that
$\mathbb{P}[{\mathcal{S}}]=\sum_{1\leq T\leq n}\mathbb{P}[\tau=T,\;{\mathcal{S}}].$
To complete the proof of estimate (21), it thus suffices to show that for any $1\leq T\leq n$ we have
$\{\tau=T,\;{\mathcal{S}}\}\subseteq\{X_{m,T}\geq 1\}.$ (22)
To justify this inclusion, we define
$I_{i}:=\mathcal{A}_{n}(G_{i}^{(\tau)})\qquad\text{for all $1\leq i\leq m$.}$ (23)
On the event $\{\tau=T,\;{\mathcal{S}}\}$, note that each $I_{i}$ is an independent set in $G_{i}^{(T)}$, and that ${\mathcal{S}}$ implies
$\min_{1\leq i\leq m}|I_{i}|=\min_{1\leq i\leq m}\bigl{|}\mathcal{A}(G_{i}^{(T)})\bigr{|}\geq(1+\varepsilon)\log_{b}(np),$
so that $(I_{1},\dots,I_{m})$ satisfies property (i) in Definition 3.7. Observing that the algorithm (see Definition 2.1) satisfies $\mathcal{A}_{n}(G_{1}^{(T)})\cap V_{T}=\mathcal{A}_{T}(G_{1}^{(T)})$, on the event $\{\tau=T\}$ it follows using the definitions (9) and (12) of $G_{1}^{(T)}=G_{1}$ and $\tau$ that
$|I_{1}\cap V_{T}|=\bigl{|}\mathcal{A}_{T}(G_{1}^{(T)})\bigr{|}=|\mathcal{A}_{T}(G_{1})|=\left\lceil(1-\varepsilon/2)\log_{b}(np)\right\rceil=N.$
Since similarly $\mathcal{A}_{n}(G_{i}^{(T)})\cap V_{T}=\mathcal{A}_{T}(G_{i}^{(T)})$, on the event $\{\tau=T\}$ it follows using Remark 3.4 that
$I_{i}\cap V_{T}=\mathcal{A}_{T}(G_{i}^{(T)})=\mathcal{A}_{T}(G_{1}^{(T)})=I_{1}\cap V_{T}\qquad\text{for all $1\leq i\leq m$,}$
so that $(I_{1},\dots,I_{m})$ satisfies property (ii) in Definition 3.7. Finally, using $E_{T}\subseteq E_{n}$, the $c$-good algorithm ${\mathcal{A}}$ ensures by Definition 2.2 that
$\bigl{|}\tbinom{I_{i}}{2}\cap E_{T}\bigr{|}\leq\bigl{|}\tbinom{\mathcal{A}_{n}(G_{i}^{(\tau)})}{2}\cap E_{n}\bigr{|}\leq c\log^{2}_{b}(np),$
so that $(I_{1},\dots,I_{m})$ satisfies property (iii) in Definition 3.7. To sum up, on the event $\{\tau=T,\;{\mathcal{S}}\}$ we showed that $(I_{1},\dots,I_{m})$ satisfies properties (i)(iii) in Definition 3.7, which by definition of $X_{m,T}$ establishes the desired inclusion (22) and thus completes the proof of Lemma 3.8 (as discussed). ∎
The following proposition intuitively states that the right-hand side of (21) and thus the success probability $\mathbb{P}[{\mathcal{S}}]$ is very small, provided that $d$ and $n$ are sufficiently large. We defer the proof of Proposition 3.9 to Section 4, since it is rather tangential to the main argument here.
Proposition 3.9.
There exist constants $d,n_{1}>1$ (which may depend on $\varepsilon,c,m$) such that, for all $n\geq n_{1}$, any $c$-good algorithm $\mathcal{A}$ satisfies
$\sum_{1\leq T\leq n}\mathbb{P}[X_{m,T}\geq 1]\leq(np)^{-\log_{b}(np)/2}$ (24)
for all edge-probabilities $p=p(n)$ satisfying $d/n\leq p\leq 1-n^{-1/d}$.
3.4 Completing the proof of Theorem 2.3
Assuming Proposition 3.9, we are now ready to complete the proof of Theorem 2.3. Indeed, combining estimate (14) from Lemma 3.6 with estimates (21) and (24) from Lemma 3.8 and Proposition 3.9, using the definitions (4) of $m=\lceil 16/\varepsilon^{2}\rceil$ and $\xi=\varepsilon^{2}/64$ it follows (for sufficiently large $d$ and $n$) that, say,
$\mathbb{P}[{\mathcal{E}}]\leq\mathbb{P}\bigl{[}{\mathcal{S}}\bigr{]}^{1/m}\leq(np)^{-\log_{b}(np)/(2m)}\leq(np)^{-\xi\log_{b}(np)}.$ (25)
Since ${\mathcal{E}}$ denotes the event that ${\bigl{|}{\mathcal{A}}(G_{1})\bigr{|}}\geq{(1+\varepsilon)\log_{b}(np)}$, using that $G_{1}\sim\mathbb{G}(n,p)$ by (7) we also have
$\mathbb{P}[{\mathcal{E}}]=\mathbb{P}_{G\sim\mathbb{G}(n,p)}\bigl{[}|\mathcal{A}(G)|\geq(1+\varepsilon)\log_{b}(np)\bigr{]}.$ (26)
Combining (25)–(26) with (6) therefore establishes the desired probability bound (2). To complete the proof of Theorem 2.3, it remains to prove Proposition 3.9, which is the subject of the next section.
4 Unlikely independent sets: proof of Proposition 3.9
In this section we give the deferred proof of Proposition 3.9, which shows that a certain family of independent sets occurs with very small probability. To avoid clutter, we set
$L:=\log_{b}(np),\qquad\ell:=\log(np),\qquad\gamma:=1-\varepsilon/2,\quad\text{and }\quad N:=\left\lceil\gamma L\right\rceil,$ (27)
and henceforth always tacitly assume that $n$ is sufficiently large whenever necessary (how large may depend on $\varepsilon,c,m,d$).
We claim that, to complete the proof of Proposition 3.9, it suffices to show that there is a constant ${d_{0}>1}$ (which may depend on $\varepsilon,c,m$) such that, for all $d\geq d_{0}$, we have
$\max_{1\leq T\leq n}\mathbb{P}[Y_{m,T}\geq 1]$ $\;\leq\;m\cdot(np)^{-L},$ (28)
$\max_{1\leq T\leq n}\mathbb{P}[Z_{m,T}\geq 1]$ $\;\leq\;n^{m}\cdot(np)^{-L},$ (29)
for all edge-probabilities $p=p(n)$ satisfying $d/n\leq p\leq 1-n^{-1/d}$, where $Y_{m,T}$ and $Z_{m,T}$ count the number of $m$-tuples $(I_{1},\dots,I_{m})$ in Definition 3.7 with ${\max_{1\leq i\leq m}|I_{i}|}>{3L}$ and ${\max_{1\leq i\leq m}|I_{i}|}\leq{3L}$, respectively. To see this claim, note that by combining $X_{m,T}=Y_{m,T}+Z_{m,T}$ with estimates (28)–(29), using the definition (27) of $L=\log_{b}(np)$ it follows (for sufficiently large $d$ and $n$) that, say,
$\begin{split}\sum_{1\leq T\leq n}\mathbb{P}[X_{m,T}\geq 1]&\leq\sum_{1\leq T\leq n}\Bigl{(}\mathbb{P}[Y_{m,T}\geq 1]+\mathbb{P}[Z_{m,T}\geq 1]\Bigr{)}\\ &\leq n\cdot 2mn^{m}\cdot(np)^{-\log_{b}(np)}.\end{split}$ (30)
Using the lower bound (5) for $\log_{b}(np)$, by distinguishing the cases $p\geq n^{-1/2}$ and $d/n\leq p\leq n^{-1/2}$, say, it is routine to see that for sufficiently large $d$ and $n$ (depending on $m,c_{0}$) we have
$(np)^{-\log_{b}(np)/2}\leq(np)^{-c_{0}\log(d)/(2p)}\leq\frac{1}{2mn^{m+1}},$ (31)
which together with (30) readily proves estimate (24) and thus Proposition 3.9, as claimed.
The remainder of this section is devoted to the proof of the probability estimates (28) and (29), for which we shall combine random graph and optimization arguments.
4.1 Proof of estimate (28): large independent sets
Fix $1\leq T\leq n$. Recall that $Y_{m,T}$ counts the number of $m$-tuples $(I_{1},\dots,I_{m})$ in Definition 3.7 with ${\max_{1\leq i\leq m}|I_{i}|}>{3L}$. Note that $Y_{m,T}\geq 1$ implies existence of $1\leq i\leq m$ such that $G_{i}^{(T)}$ contains an independent set $I$ of size
$k:=\left\lceil 3L\right\rceil,$
where $G_{i}^{(T)}\sim\mathbb{G}(n,p)$ by Remark 3.3. Using estimate (5) and $L=\log_{b}(np)$, note that for sufficiently large $d$ we have $ne/k\leq np$ and $(k-1)/2\geq 4L/3$, say, so that a standard union bound argument (over all possible choices of random graphs $G_{i}^{(T)}$ and independent sets $I$) and the well-known binomial coefficient estimate $\binom{n}{k}\leq(ne/k)^{k}$ yield
$\begin{split}\mathbb{P}[Y_{m,T}\geq 1]&\leq\sum_{1\leq i\leq m}\sum_{I\subseteq[n]:|I|=k}(1-p)^{\binom{k}{2}}\\ &\leq m\cdot\binom{n}{k}(1-p)^{\binom{k}{2}}\leq m\cdot\biggl{(}\frac{ne}{k}\cdot(1-p)^{(k-1)/2}\biggr{)}^{k}\\ &\leq m\cdot\Bigl{(}np\cdot(np)^{-4/3}\Bigr{)}^{k}\leq m\cdot(np)^{-L},\end{split}$ (32)
completing the proof of estimate (28).
4.2 Proof of estimate (29): overlapping independent sets
Fix $1\leq T\leq n$. Recall that $Z_{m,T}$ counts the number of $m$-tuples $(I_{1},\dots,I_{m})$ in Definition 3.7 with ${\max_{1\leq i\leq m}|I_{i}|}\leq{3L}$. Given $\boldsymbol{\alpha}=(\alpha_{1},\dots,\alpha_{m})\in[1+\varepsilon,3]^{m}$, we introduce the auxiliary random variable $Z_{\boldsymbol{\alpha},m,T}$ that counts the number of all $(I_{1},\dots,I_{m})$ arising in Definition 3.7 with the additional constraint that $|I_{i}|=\alpha_{i}L$ for all $i\in[m]$. Note that $\boldsymbol{\alpha}L\in\mathbb{N}^{m}$ necessarily, and that in fact $\boldsymbol{\alpha}L\in[n]^{m}$ holds. Hence
$Z_{m,T}=\sum_{\begin{subarray}{c}\boldsymbol{\alpha}\in[1+\varepsilon,3]^{m}:\\ \boldsymbol{\alpha}L\in[n]^{m}\end{subarray}}Z_{\boldsymbol{\alpha},m,T}.$ (33)
Using Markov’s inequality and linearity of expectation, we thus infer that
$\mathbb{P}[Z_{m,T}\geq 1]\leq\mathbb{E}[Z_{m,T}]=\sum_{\begin{subarray}{c}\boldsymbol{\alpha}\in[1+\varepsilon,3]^{m}:\\ \boldsymbol{\alpha}L\in[n]^{m}\end{subarray}}\mathbb{E}[Z_{\boldsymbol{\alpha},m,T}].$ (34)
We shall bound $\mathbb{E}[Z_{\boldsymbol{\alpha},m,T}]$ in the following subsections, using counting and probability estimates.
4.2.1 Counting estimates
We start with the following counting estimate for the number of $m$-tuples $(I_{1},\dots,I_{m})$ with specific sizes. Recall that $V_{T}=\{v_{1},\ldots,v_{T}\}$ denotes the first $T$ vertices considered by the algorithm ${\mathcal{A}}$, which we assume to be known in Lemma 4.1 below.
Lemma 4.1.
For sufficiently large $d$ the following is true for any $\boldsymbol{\alpha}=(\alpha_{1},\dots,\alpha_{m})\in[1+\varepsilon,3]^{m}$. Given $V_{T}\subseteq[n]$, the total number of $(I_{1},\dots,I_{m})$ with $I_{1},\dots,I_{m}\subseteq[n]$ that satisfy property (ii) in Definition 3.7 and $|I_{i}|=\alpha_{i}L$ for all $1\leq i\leq m$ is at most
$\exp\biggl{(}\ell L\cdot\Bigl{[}\sum_{1\leq i\leq m}\bigl{(}\alpha_{i}-\gamma\bigr{)}+\gamma\Bigr{]}\biggr{)}.$ (35)
Proof.
To bound the total number of $(I_{1},\dots,I_{m})$, we will describe a way to construct all such tuples (and potentially more tuples), and bound the number of choices in our construction from above. We first choose the set $I_{1}\cap V_{T}$ of ${N=\left\lceil\gamma L\right\rceil}$ common vertices, which using $V_{T}\subseteq[n]$ can be done in at most
$\binom{|V_{T}|}{N}\leq\binom{n}{\left\lceil\gamma L\right\rceil}$
many ways. Given $I_{1}\cap V_{T}$ of size $|I_{1}\cap V_{T}|=N$, note that property (ii) in Definition 3.7 enforces $I_{i}\cap V_{T}=I_{1}\cap V_{T}$ for each $1\leq i\leq m$. With this in mind, we next choose the remaining vertices of the sets $I_{i}$ of size $|I_{i}|=\alpha_{i}L$, which using $|I_{1}\cap V_{T}|=N=\left\lceil\gamma L\right\rceil$ can be done in at most
$\prod_{1\leq i\leq m}\binom{n}{\alpha_{i}L-\left\lceil\gamma L\right\rceil}$
many ways. Using the standard estimate $\binom{n}{k}\leq(ne/k)^{k}$, it follows that the total number of $(I_{1},\dots,I_{m})$ that satisfy property (ii) in Definition 3.7 and $|I_{i}|=\alpha_{i}L$ for all $1\leq i\leq m$ is at most
$\binom{n}{\left\lceil\gamma L\right\rceil}\cdot\prod_{1\leq i\leq m}\binom{n}{\alpha_{i}L-\left\lceil\gamma L\right\rceil}\leq\biggl{(}\frac{ne}{\left\lceil\gamma L\right\rceil}\biggr{)}^{\left\lceil\gamma L\right\rceil}\cdot\prod_{1\leq i\leq m}\biggl{(}\frac{ne}{\alpha_{i}L-\left\lceil\gamma L\right\rceil}\biggr{)}^{\alpha_{i}L-\left\lceil\gamma L\right\rceil},$
[…]
$\begin{split}\bigl{(}np\bigr{)}^{\left\lceil\gamma L\right\rceil}\cdot\prod_{1\leq i\leq m}\bigl{(}np\bigr{)}^{\alpha_{i}L-\left\lceil\gamma L\right\rceil}&=\bigl{(}np\bigr{)}^{\sum_{1\leq i\leq m}\alpha_{i}L-(m-1)\left\lceil\gamma L\right\rceil},\end{split}$
which using $np=e^{\ell}$ readily completes the proof of Lemma 4.1. ∎
4.2.2 Probability estimates
We next prove the following probability estimate. Similar as in the proof of Lemma 3.6, we henceforth let $\Xi_{T}$ encode the edge-status of all pairs $\{u,v\}\in\tbinom{V_{T}}{2}\cup E_{T}$, i.e., the edge-status of all pairs revealed during the first $T$ rounds of the algorithm ${\mathcal{A}}$. Recalling that the algorithm ${\mathcal{A}}$ is deterministic, note that $V_{T}=\{v_{1},\ldots,v_{T}\}$ and $E_{T}=S_{1}\cup\cdots\cup S_{T}$ are both determined by $\Xi_{T}$, where $E_{T}$ contains the ‘exceptional’ pairs revealed in (b) during the first $T$ rounds of ${\mathcal{A}}$.
Lemma 4.2.
For sufficiently large $d$ the following is true for any $\boldsymbol{\alpha}=(\alpha_{1},\dots,\alpha_{m})\in[1+\varepsilon,3]^{m}$. For every $m$-tuple $(I_{1},\dots,I_{m})$ with $I_{1},\dots,I_{m}\subseteq[n]$ that satisfies properties (ii) and (iii) in Definition 3.7 as well as $|I_{i}|=\alpha_{i}L$ for all $i\in[m]$, we have
$\begin{split}&\mathbb{P}\Bigl{[}\text{$I_{i}$ is an independent set in $G_{i}^{(T)}$ for all $1\leq i\leq m$}\>\Big{|}\>\Xi_{T}\Bigr{]}\\ &\quad\leq\exp\biggl{(}-\ell L\cdot\sum_{1\leq i\leq m}\Bigl{(}\alpha_{i}^{2}/2-\gamma^{2}/2-2c\Bigr{)}\biggr{)}.\end{split}$ (36)
Proof.
Given a set $I_{i}\subseteq[n]$ as in the statement of the lemma, we define the set
$\begin{split}F_{i}&:=\tbinom{I_{i}}{2}\setminus\Bigl{(}\tbinom{V_{T}}{2}\cup E_{T}\Bigr{)},\end{split}$ (37)
which is determined by $\Xi_{T}$ (since $V_{T}$ and $E_{T}$ are both determined by $\Xi_{T}$). Focusing on the absence of the edges in the $F_{i}$ (ignoring the absence of other potential edges might seem wasteful, but later on we can compensate for this in (41)–(43) by taking $m$ sufficiently large) it follows that
$\mathbb{P}\Bigl{[}\text{$I_{i}$ is an independent set in $G_{i}^{(T)}$ for all $1\leq i\leq m$}\>|\>\Xi_{T}\Bigr{]}$ $\leq\mathbb{P}[{\mathcal{F}}\>|\>\Xi_{T}]$ (38)
for the ‘forbidden edges’ event
:=$\begin{split}{\mathcal{F}}:=\Bigl{\{}e_{i,T}\bigl{(}\{u,v\}\bigr{)}=0\text{ for all $\{u,v\}\in F_{i}$ and $i\in[m]$}\Bigr{\}},\end{split}$ (39)
where $e_{i,T}\bigl{(}\{u,v\}\bigr{)}\in\{0,1\}$ denotes the edge-status of the pair $\{u,v\}$ in $G_{i}^{(T)}$, see Section 3.2. We next give lower bounds on the sizes of the sets $F_{i}$ defined in (37), exploiting that properties (ii) and (iii) in Definition 3.7 ensure that $|I_{i}\cap V_{T}|=|I_{1}\cap V_{T}|=N=\left\lceil\gamma L\right\rceil$ and $\bigl{|}\binom{I_{i}}{2}\cap E_{T}\bigr{|}\leq c\log_{b}^{2}(np)=cL^{2}$. Namely, together with $|I_{i}|=\alpha_{i}L$ and $\alpha_{i}+\gamma\leq 4$, it follows that we deterministically have, say,
$\begin{split}|F_{i}|&\geq\bigl{|}\tbinom{I_{i}}{2}\bigr{|}-\bigl{|}\tbinom{I_{i}\cap V_{T}}{2}\bigr{|}-\bigl{|}\tbinom{I_{i}}{2}\cap E_{T}\bigr{|}\\ &\geq(\alpha_{i}L-1)^{2}/2-(\gamma L+1)^{2}/2-cL^{2}\\ &=L^{2}\Bigl{(}\alpha_{i}^{2}/2-\gamma^{2}/2-(\alpha_{i}+\gamma)/L-c\Bigr{)}\\ &\geq L^{2}\Bigl{(}\alpha_{i}^{2}/2-\gamma^{2}/2-4/L-c\Bigr{)}.\end{split}$ (40)
Note that all pairs $\{u,v\}\in F_{1}\cup\cdots\cup F_{m}$ satisfy $\{u,v\}\not\in\tbinom{V_{T}}{2}\cup E_{T}$, which means that they have not been revealed during the first $T$ rounds of the algorithm. By inspecting (11) it therefore follows that, conditional on $\Xi_{T}$, all the pairs in $F_{1}\cup\cdots\cup F_{m}$ are included as edges into each graph $G_{1}^{(T)},\ldots,G_{m}^{(T)}$ independently with probability $p$. In view of (39) and the fact that $F_{i}$ is determined by $\Xi_{T}$, using the deterministic lower bound (40) on $|F_{i}|$ and $L=\log_{b}(np)=-\ell/\log(1-p)$, it thus follows that
$\begin{split}\mathbb{P}[{\mathcal{F}}\>|\>\Xi_{T}]&=\prod_{1\leq i\leq m}(1-p)^{|F_{i}|}\\ &\leq\prod_{1\leq i\leq m}(1-p)^{L^{2}\bigl{(}\alpha_{i}^{2}/2-\gamma^{2}/2-4/L-c\bigr{)}}\\ &=\exp\biggl{(}-\ell L\cdot\sum_{1\leq i\leq m}\Bigl{(}\alpha_{i}^{2}/2-\gamma^{2}/2-4/L-c\Bigr{)}\biggr{)}.\end{split}$
Recalling that $L=\log_{b}(np)\geq c_{0}\log d$ by (5), for sufficiently large $d$ (depending on $c$) we infer that ${4/L\leq c}$, which together with (38) completes the proof of estimate (36) and thus Lemma 4.2. ∎
4.2.3 Completing the proof: putting everything together
After these preparations we are now ready to bound $\mathbb{E}[Z_{\boldsymbol{\alpha},m,T}]$, by combining the number of possible $(I_{1},\dots,I_{m})$ via the counting estimate from Lemma 4.1, with the probability estimate from Lemma 4.2. Indeed, since the vertex-set $V_{T}$ is determined by $\Xi_{T}$ (as mentioned above Lemma 4.2), by multiplying estimates (35) and (36) it readily follows for any $\boldsymbol{\alpha}=(\alpha_{1},\dots,\alpha_{m})\in[1+\varepsilon,3]^{m}$ with $\boldsymbol{\alpha}L\in[n]^{m}$ that
$\mathbb{E}[Z_{\boldsymbol{\alpha},m,T}]=\mathbb{E}\bigl{[}\mathbb{E}[Z_{\boldsymbol{\alpha},m,T}\>|\>\Xi_{T}]\bigr{]}\leq\mathbb{E}\Bigl{[}\exp\Bigl{(}-\ell L\cdot\Psi(\boldsymbol{\alpha})\Bigr{)}\Bigr{]}=\exp\Bigl{(}-\ell L\cdot\Psi(\boldsymbol{\alpha})\Bigr{)}$ (41)
for the function
$\Psi(\boldsymbol{\alpha}):=\sum_{1\leq i\leq m}\Bigl{[}\bigl{(}\alpha_{i}^{2}/2-\alpha_{i}\bigr{)}-\bigl{(}\gamma^{2}/2-\gamma\bigr{)}-2c\Bigr{]}-\gamma.$ (42)
We now bound $\Psi(\boldsymbol{\alpha})$ from below. Recalling that $\gamma=1-\varepsilon/2$, we have
$\gamma^{2}/2-\gamma=\varepsilon^{2}/8-1/2.$
Furthermore, by noting that the following minimum is attained for $\alpha_{i}=1+\varepsilon$, we have
$\min_{\alpha_{i}\in[1+\varepsilon,3]}\bigl{(}\alpha_{i}^{2}/2-\alpha_{i}\bigr{)}=\varepsilon^{2}/2-1/2.$
Inserting these two estimates into (42), we now exploit that by definition (4) we have $42c\leq\varepsilon^{2}/$ 4 and $m\geq 16/\varepsilon^{2}$, so that using $\gamma\leq 1$ we infer that
$\begin{split}\Psi(\boldsymbol{\alpha})&\geq m\cdot\bigl{(}\varepsilon^{2}/2-\varepsilon^{2}/8-\varepsilon^{2}/4\bigr{)}-\gamma\geq m\cdot\varepsilon^{2}/8-1\geq 1.\end{split}$ (43)
To sum up, since there are at most $n^{m}$ choices of $\boldsymbol{\alpha}\in[1+\varepsilon,3]^{m}$ for which $\boldsymbol{\alpha}L\in[n]^{m}$, by inserting (41) and (43) into (34) it follows that
$\mathbb{P}[Z_{m,T}\geq 1]\leq\sum_{\begin{subarray}{c}\boldsymbol{\alpha}\in[1+\varepsilon,3]^{m}:\\ \boldsymbol{\alpha}L\in[n]^{m}\end{subarray}}\mathbb{E}[Z_{\boldsymbol{\alpha},m,T}]\leq n^{m}\cdot\exp\bigl{(}-\ell L\bigr{)},$
which using $\ell=\log(np)$ completes the proof of estimate (29) and thus Proposition 3.9, as discussed.
5 Achievability result: proof of Theorem 2.4
In this section we prove our achievability result Theorem 2.4, which shows that a certain online algorithm can find an independent set of size $(1+\varepsilon)\log_{b}(n)$ with high probability. Heuristically speaking, the algorithm we shall use consists of three phases:
  1. (i)
    first we greedily find an independent set $I_{T}$ of size $|I_{T}|\approx(1-\varepsilon)\log_{b}(n)$,
  2. (ii)
    then we iteratively find a vertex-set $R$ of size $|R|\approx n^{\varepsilon}$ that only contains vertices $w\not\in I_{T}$ which have no neighbors in $I_{T}$, and
  3. (iii)
    finally we find using brute-force search an independent set $J\subseteq R$ of size approximately $|J|\approx 2\log_{b}(|R|)\approx 2\varepsilon\log_{b}(n)$, and construct an independent set $I:=I_{T}\cup J$ of size $|I|\approx(1+\varepsilon)\log_{b}(n)$.
Turning to the details, we now describe a deterministic algorithm ${\mathcal{A}}$ that implements the above-mentioned three-phase idea in a way that fits it into the online-framework of Definition 2.1. Setting
$\ell:=\bigl{\lfloor}(1-\varepsilon)\log_{b}(n)-3.5\log_{b}\log_{b}(n)\bigr{\rfloor},\qquad T:=\ell\cdot\left\lceil n^{1-\varepsilon}\ell\right\rceil,\quad\text{ and }\quad r:=\left\lfloor n^{\varepsilon}\log^{3}_{b}(n)\right\rfloor,$ (44)
the algorithm ${\mathcal{A}}$ initially starts with $I_{0}:=\emptyset$, and then proceeds in rounds as follows.
  • Greedy rounds: In round $1\leq t\leq T$ we consider vertex $t\in[n]$, and reveal the edge-status of all pairs $\{u,t\}$ with $u\in I_{t-1}$. If either $|I_{t-1}|\geq\ell$ or vertex $t$ has a neighbor in $I_{t-1}$, then we set $I_{t}:=I_{t-1}$. Otherwise, we set $I_{t}:=I_{t-1}\cup\{t\}$.
  • Search round: In round $t=T+1$, we consider vertex $n$ and set $I_{T+1}:=I_{T}$. In this round we use the exceptional set $S_{T+1}$ to reveal the edge-status of all pairs $\{u,v\}$ with $u\in I_{T}$ and $v\in\{T+1,\ldots,n-2\}=:W$, in order to determine the set $R\subseteq W$ of vertices in $W$ that have no neighbors in $I_{T}$. If $|R|>r$, then we iteratively remove vertices from $R$ (in lexicographic order, say) until the resulting set satisfies $|R|=r$.
  • Brute-force round: In round $t=T+2$, we consider vertex $n-1$ and set $I_{T+2}:=I_{T+1}$. In this round we use the exceptional set $S_{T+2}$ to reveal the edge-status of all pairs ${\{u,v\}\in\tbinom{R}{2}}$, and then use brute-force search to find a largest independent set $J\subseteq R$ of size at most ${|J|\leq\left\lfloor 2\log_{b}(r)\right\rfloor}$, where we write ${J=\{w_{1},\ldots,w_{|J|}\}}$.
  • Final rounds: In rounds $T+3\leq t\leq T+2+|J|$, we consider vertex $w_{t-(T+2)}\in[n]$ and set $I_{t}:=I_{t-1}\cup\{w_{t-(T+2)}\}$. In the remaining rounds $T+3+|J|\leq t\leq n$ we sequentially consider the remaining vertices in $W\setminus J$, and each time set $I_{t}:=I_{t-1}$.
Note that algorithm ${\mathcal{A}}$ has running time $n^{\Theta(1)}+n^{\Theta(1)}\cdot\tbinom{n}{\left\lfloor 2\log_{b}(r)\right\rfloor}=n^{(2+o(1))\varepsilon\log_{b}(n)}$, where here and below we always tacitly assume that $n$ is sufficiently large whenever necessary (how large may depend on ${\varepsilon,p}$). Furthermore, we have $I_{n}=I_{T}\cup J$ and $E_{n}={S_{1}\cup\cdots\cup S_{n}}={S_{T+1}\cup S_{T+2}}$, so using $|J|\leq\left\lfloor 2\log_{b}(r)\right\rfloor$ and $|I_{T}|\leq\ell$ the algorithm ${\mathcal{A}}$ deterministically satisfies, say,
$\begin{split}\bigl{|}\tbinom{I_{n}}{2}\cap E_{n}\bigr{|}&\leq|J|\cdot|I_{T}|+\tbinom{|J|}{2}\\ &\leq 2\log_{b}(r)\cdot\bigl{[}\ell+\log_{b}(r)\bigr{]}\\ &\leq\bigl{[}2\varepsilon\log_{b}(n)+6\log_{b}\log_{b}(n)\bigr{]}\cdot\bigl{[}\log_{b}(n)-0.5\log_{b}\log_{b}(n)\bigr{]}\\ &\leq 3\varepsilon\log^{2}_{b}(np),\end{split}$ (45)
so that ${\mathcal{A}}$ is $3\varepsilon$-good according to Definition 2.2. Note that the algorithm ${\mathcal{A}}$ finds an independent set of size $|I_{n}|=|I_{T}|+|J|$. To complete the proof of Theorem 2.4, it thus suffices to prove the following lemma.
Lemma 5.1.
We have $\mathbb{P}[|I_{T}|+|J|\geq(1+\varepsilon)\log_{b}(n)]\to 1$ as $n\to\infty$.
The proof of Lemma 5.1 uses the following auxiliary result, whose proof we defer.
Lemma 5.2.
We have $\mathbb{P}[|I_{T}|=\ell,\>|R|=r]\to 1$ as $n\to\infty$.
Proof of Lemma 5.1.
We henceforth condition on the outcome of the greedy rounds and the search round, where by Lemma 5.2 we may assume that $|I_{T}|=\ell$ and $|R|=r$. Since all pairs in $\binom{R}{2}$ are included as edges independently with probability $p$ (since their status has not yet been revealed by the algorithm), it follows that $G[R]\sim\mathbb{G}(r,p)$. Using well-known random graph results (see [JLR00, Theorem 7.1], for example) on the size of the largest independent of $\mathbb{G}(r,p)$, as $n\to\infty$ and thus $r\to\infty$ it follows that
$\mathbb{P}\bigl{[}\alpha(G[R])\geq 2\log_{b}(r)-2\log_{b}\log_{b}(rp)+2\log_{b}(e/2)+1-1/p\bigr{]}\to 1,$
where we omitted the conditioning from our notation to avoid clutter. Using $r=\left\lfloor n^{\varepsilon}\log^{3}_{b}(n)\right\rfloor$ and the definition (44) of $\ell$, with probability tending to one as $n\to\infty$, we thus have
$\begin{split}|I_{T}|+|J|&=\ell+\min\bigl{\{}\alpha(G[R]),\>\left\lfloor 2\log_{b}(r)\right\rfloor\bigr{\}}\\ &\geq\bigl{[}(1-\varepsilon)\log_{b}(n)-3.5\log_{b}\log_{b}(n)-1\bigr{]}+\bigl{[}2\varepsilon\log(n)+4\log_{b}\log_{b}(n)-O(1)\bigr{]}\\ &>(1+\varepsilon)\log_{b}(n),\end{split}$ (46)
which completes the proof of Lemma 5.1. ∎
Finally, we give the deferred proof of the auxiliary result Lemma 5.2.
Proof of Lemma 5.2.
We start by showing that $\mathbb{P}(|I_{T}|=\ell)\to 1$, for which our argument is based on the analysis of Grimmett and McDiarmid [GM75] from 1975. Let $\Delta_{i}$ denote the number of rounds required for the independent set to grow from size $i-1$ to $i$. Note that a union bound argument gives
$\begin{split}\mathbb{P}[|I_{T}|<\ell]&=\mathbb{P}[\Delta_{1}+\cdots+\Delta_{\ell}>T]\\ &\leq\mathbb{P}[\Delta_{i}>T/\ell\text{ for some $1\leq i\leq\ell$}]\\ &\leq\sum_{1\leq i\leq\ell}\mathbb{P}[\Delta_{i}>T/\ell].\end{split}$
Since in round $t$ all pairs in $I_{t-1}\times\{t\}$ are included as edges independently with probability $p$ (since their status has not yet been revealed by the algorithm), using the standard estimate ${1-x\leq e^{-x}}$ it follows that
$\begin{split}\mathbb{P}[\Delta_{i}>T/\ell]\leq\bigl{(}1-(1-p)^{i-1}\bigr{)}^{T/\ell}\leq e^{-(1-p)^{i-1}T/\ell}.\end{split}$
Using $i-1\leq\ell$ and the definitions (44) of $\ell$ and $T$, it readily follows that
$\begin{split}\mathbb{P}[|I_{T}|<\ell]\leq\ell\cdot e^{-(1-p)^{\ell}T/\ell}\leq\ell e^{-\ell}\to 0,\end{split}$
which together with the deterministic bound $|I_{T}|\leq\ell$ (guaranteed by construction of the algorithm) shows that $\mathbb{P}(|I_{T}|=\ell)\to 1$,
We henceforth condition on the outcome of the greedy rounds, and assume that $|I_{T}|=\ell$ by the analysis above. In this conditional space we will show that $\mathbb{P}(|R|=r)\to 1$, where we here and below omit the conditioning from our notation (to avoid clutter). Let $X$ denote the number of vertices in $W$ that have no neighbors in $I_{T}$. Since all pairs in $W\times I_{T}$ are included as edges independently with probability $p$ (since their status has not yet been revealed by algorithm), we have
$X\sim{\rm Binomial}\Bigl{(}|W|,\>(1-p)^{\ell}\Bigr{)}.$
Recalling that $|W|=n-(T+2)$, by definition (44) of $\ell$ and $r$ we thus have, say,
$\mathbb{E}[X]=|W|\cdot(1-p)^{\ell}\geq(1-o(1))\cdot n^{\varepsilon}\log_{b}^{3.5}(n)>2r.$
Using standard Chernoff bounds (see [JLR00, Theorem 2.1]), it follows that
$\mathbb{P}[|R|<r]=\mathbb{P}[X<r]\leq\mathbb{P}\bigl{[}X\leq\mathbb{E}[X]/2\bigr{]}\leq e^{-\mathbb{E}[X]/8}\leq e^{-r/4}\to 0,$
which together with the deterministic bound $|R|\leq r$ (guaranteed by construction of the algorithm) shows that $\mathbb{P}[|R|=r]\to 1$, completing the proof of Lemma 5.2. ∎